rate of 5.25%, 1 Lender I offers you a fixed rate 15-year mortgage at an annual interes compounded monthly, with no points a. F ind your monthly payments under this option. b. Find the total amount of money paid to the lender. c. Find the total amount of interest you will pay over the life of the loan. 2. Lender II offers you a fixed rate 30-year mortgage at an annual interest rate of 5.75%, compounded monthly, with one point a. Find your monthly payments under this option. (Dont forget to include the point point in the loan amount.) b. Find the total amount of money paid to the lender. c. Find the total amount of interest you will pay over the life of the loan.
Answer #1

Assume you need $1000 for your mortgage in both scenarios

1) Monthly Payment can be calculated using PMT function

N = 15 x 12 = 180, I/y = 5.25%/12, PV = 1000, FV = 0 => Compute PMT = $8.04

Total Amount Paid = 8.04 x 180 = $1,446.98

Interest Paid = 1,446.98 - 1,000 = $446.98

2) N = 30 x 12 = 360, I/Y = 5.75% / 12, PV = 1000 x (1 + 1%) = 1,010, FV = 0

=> Compute PMT = $5.89

Total Amount Paid = 5.89 x 30 x 12 = $2,121.87

Interest Paid = 2,121.87 - 1,000 = 1,121.87